# Information Inequalities for Five Random Variables

### v1

@@ Abstract.
The entropic region is formed by the collection of the Shannon entropies of all subvectors of finitely many jointly distributed discrete random variables. For four or more variables the structure of the entropic region is mostly unknown. We utilize a variant of the Maximum Entropy Method to delimit the five-variable entropy region. This method adds copies of some of the random variables in generations. A significant reduction in computational complexity, achieved through theoretical considerations and by harnessing the inherent symmetries, allowed us to calculate all five-variable non-Shannon inequalities provided by the first nine generations. Based on the results, we define two infinite collections of such inequalities, and prove them to be entropy inequalities. We investigate downward closed subsets of non-negative lattice points that parameterize these collections, based on which we develop an algorithm to enumerate all extremal inequalities. The discovered set of entropy inequalities is conjectured to characterize the applied method completely.
@@ 1. Introduction
Many important mathematical problems can be reduced to the following question: does a collection of finite random variables exist such that the entropies of the variable subsets satisfy certain linear constraints? Examples include, but are not limited to, channel coding [13] and network coding in particular [36], estimating the efficiency of secret sharing schemes [3, 4, 20], questions about matroid representations [2], guessing games [32], extracting information from common strings in cryptography [17], additive combinatorics [23], and finding conditional independence inference rules [34].
@@
Previous works on generating and applying non-Shannon entropy inequalities, such as [4, 8, 9, 15, 29, 34], focused on the four-variable case, and only a few sporadic five-variable non-Shannon inequalities have been discovered, such as the MMRV inequality from [24]. This is the first work that provides a method which generates an infinite collection of non-Shannon bounds on the five-variable entropy region $\Gamma^{*}_{\!5}$. Compared to the four-variable case, there are significant challenges, both theoretical and computational. The four-variable entropy region $\Gamma^{*}_{\!4}$ sits in the $15$-dimensional Euclidean space, while the five-variable region $\Gamma^{*}_{\!5}$ is $31$-dimensional. The structure of the Shannon bound $\Gamma_{\!4}$ is well-understood: it has $41$ extremal directions, and only $6$ of them have no entropic points. The entropy region $\Gamma^{*}_{\!4}$ has an inner polyhedral cone where it fills its Shannon bound, and has six isomorphic “protrusions” towards the six exceptional extremal directions, each protrusion surrounded by $15$ hyperplanes of which $14$ come from the Shannon bound [25]. Only the protrusions contribute to new entropy inequalities, and their dimension can be reduced to $10$. Computational results about $\Gamma^{*}_{\!4}$ can be obtained by computing vertices and facets of numerous implicitly defined $10$-dimensional polyhedra [9]. In contrast, the Shannon bound $\Gamma_{\!5}$ of the five-variable entropy region has 117,983 extremal directions [33], and for a few of them it is not even known whether they contain an entropic point or not. No structural reduction similar to the four-variable case is available, and it is not known whether such a reduction exists or not. Computations about $\Gamma^{*}_{\!5}$ can still be reduced to $25$-dimensional polyhedral enumeration problems (although with significantly larger number of constraints than in the $4$-variable case). The complexity of enumeration problems typically doubles when the dimension increases by one, making such high-dimensional enumeration problems practically intractable.
@@
We overcome this computational difficulty by applying a particular variant of the Maximum Entropy Method. This variant, working in generations, first reduces the problem dimension from $31$ to $19$, and then, at each generation, adds extra copies of some of the random variables increasing the problem dimension again. Theoretical considerations and harnessing the inherent symmetry allowed us to complete the associated polyhedral computations up to nine generations. The output was the complete list of five-variable non-Shannon inequalities provided by the first nine generations. Based on the experimental results, we define an infinite collection of five-variable inequalities that we prove to be provided by this MEM variant—in particular, they are valid non-Shannon entropy inequalities—, and conjecture this collection to be complete, that is, no additional inequalities are yielded by this MEM variant. The collection of the inequalities is parametrized by finite, downward closed subsets of the non-negative lattice points of the plane. Some of the inequalities in our collection are consequences of the others; those that are not, are called extremal. We developed an incremental algorithm that enumerates, from generation to generation, the parameters yielding the extremal inequalities, in complete agreement with the computational results. Finally, we have looked at the large-scale behavior of the extremal inequalities, and pictured how these inequalities delimit a $3$-dimensional cross-section of $\Gamma^{*}_{\!5}$.
@@
The remaining part of the paper is organized as follows. Notations are recalled in Section 2. Section 3 describes the special variant of the Maximum Entropy Method we apply to $\Gamma^{*}_{\!5}$. Section 4 discusses possible simplifications, including how symmetry can be utilized, and how the MEM parameters were chosen. Section 5 describes the chosen coordinate systems, polyhedral computations and their results. Section 6 presents the five-variable inequalities we obtained, paving the way for the definition of two infinite families of such inequalities in Section 7. Additional theoretical results, including the proof that inequalities in these families are indeed generated by the MEM method, are presented in Section 7. Section 8 discusses methods that can recognize extremal inequalities, the incremental algorithm that enumerates the extremal inequalities for each MEM generation, describes the large-scale behavior of the new inequalities, and investigates the delimited part of the five-variable entropy region. Finally, Section 9 summarizes our work, lists open questions, and provides directions for further work.
@@ 2. Preliminaries
A discrete random variable $\xi$ takes its values from a finite set $\mathcal{X}$, called alphabet. The probability that $\xi$ takes $x\in\mathcal{X}$ is denoted by $\Pr(\xi=x)$, or simply by $\Pr(x)$ when the random variable $\xi$ is clear from the context. Suppose $\xi$ is defined on the direct product $\mathcal{X}=\prod_{i\in N}\mathcal{X}_{i}$ for some finite set $N$, called the base set. For a non-empty $A\subseteq N$ the marginal $\xi_{A}$ is defined on the product alphabet $\mathcal{X}_{A}=\prod_{i\in A}\mathcal{X}_{i}$ so that the probability of $y\in\mathcal{X}_{A}$ is the sum of the probability of those $x\in\mathcal{X}$ whose projection to $\mathcal{X}_{A}$ equals $y$:
@@
with the convention that $0\log 0=0$. If $\xi=(\xi_{i}:i\in N)$ is a joint distribution, then we write $\mathbf{H}_{\xi}(A)$ for $\mathbf{H}(\xi_{A})$. The index $\xi$ is also dropped when it is clear from the context. By convention, $\mathbf{H}(\emptyset)=0$. The entropies $\mathbf{H}_{\xi}(A)$ are arranged into a vector indexed by the non-empty subsets $A$ of $N$. This vector is the entropy profile of the distribution $\xi$. The collection of these $(2^{|N|}\mkern 1.5mu{-}\mkern 1.5mu1)$-dimensional vectors forms the entropy region $\Gamma^{*}_{\!N}$. Elements of $\Gamma^{*}_{\!N}$ are considered interchangeably as vectors, as points in this Euclidean space, and as functions assigning non-negative real numbers to non-empty subsets of the base set $N$.
@@
$f(A\mkern 1.0mu{|}\mkern 1.0muB)$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}f(AB)-f(B),$
@@
$f(A,B\mkern 1.0mu{|}\mkern 1.0muC)$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}f(AC)+f(BC)-f(ABC)-f(C),\mbox{ and}$
$f[A,B,C,D]$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-f(A,B)+f(A,B\mkern 1.0mu{|}\mkern 1.0muC)+f(A,B\mkern 1.0mu{|}\mkern 1.0muD)+f(C,D).$
The first three expressions are called conditional entropy, mutual information, and conditional mutual information, respectively. The last line defines the Ingleton expression. An entropy function is not defined on the empty set, nevertheless, $f(\emptyset)=0$ will be assumed whenever convenient. In particular, $f(A,B\mkern 1.0mu{|}\mkern 1.0mu\emptyset)$ and $f(A,B)$ are the same expressions. Frequently, when clear from the context, the function $f$ is omitted before the parenthesized expression. Additionally, if applied to singletons, the Ingleton expression is written without commas. An example is the inequality
$abcd\mkern 1.5mu]+(a,b\mkern 1.0mu{|}\mkern 1.0muz)+(b,z\mkern 1.0mu{|}\mkern 1.0mua)+(a,z\mkern 1.0mu{|}\mkern 1.0mub)+3(z\mkern 1.0mu{|}\mkern 1.0muab)\geqslant 0.$ (1)
Shannon inequalities state the non-negativity of the conditional entropy, mutual information, and conditional mutual information for all subsets $A$, $B$, $C$ of the base set $N$. They are consequences of the unique minimal set of such inequalities, called basic Shannon inequalities, see [35], listed in (B1) and (B2) below:
@@
    $f(i\mkern 1.0mu{|}\mkern 1.0muN\mkern 1.5mu{-}\mkern 1.5mui)\geqslant 0$ for all $i\in N$;
@@
    $f(a,b\mkern 1.0mu{|}\mkern 1.0muK)\geqslant 0$ for all $K\subseteq N$ and different $a,b\in N\mkern 1.5mu{-}\mkern 1.5muK$, including $K=\emptyset$.
The collection of all $(2^{|N|}\mkern 1.5mu{-}\mkern 1.5mu1)$-dimensional vectors (or points, or functions) that satisfy the Shannon inequalities is denoted by $\Gamma_{\!N}$. It is a natural outer bound for the entropy region $\Gamma^{*}_{\!N}$. $\Gamma_{\!N}$ is a pointed polyhedral cone [38]; its facets are the hyperplanes specified by the basic Shannon inequalities. Polymatroids are elements of $\Gamma_{\!N}$ written in functional form. A polymatroid is usually written as $(f,N)$, or just $f$, and we say that $f$ is on $N$. The polymatroid $f$ is entropic if it is in $\Gamma^{*}_{\!N}$, and almost entropic, or aent for short, if it is in the closure (in the usual Euclidean topology) of $\Gamma^{*}_{\!N}$. Linear inequalities valid for all polymatroids are consequences of the basic Shannon inequalities, an example is the inequality (1). A non-Shannon inequality is a homogeneous linear inequality that is valid for points of the entropic region but not for all points of the Shannon bound. Equivalently, the non-negative side of the hyperplane corresponding to such an inequality contains the complete entropy region, while it cuts properly into $\Gamma_{\!N}$.
@@
In this case we also have $f(Y_{I})=\sum_{i\in I}f(Y_{i})$ for every subset $I$ of the indices. The disjoint subsets $Y_{1}$ and $Y_{2}$ are conditionally independent over $X$ if $f(Y_{1},Y_{2}\mkern 1.0mu{|}\mkern 1.0muX)=0$; and $Y_{1},\dots,Y_{k}$ are completely conditionally independent over $X$ if $Y_{I}$ and $Y_{J}$ are conditionally independent over $X$ for arbitrary disjoint subsets $I$ and $J$ of the indices. An equivalent condition is
$f(Y_{1}\cdots Y_{k}\mkern 1.0mu{|}\mkern 1.0muX)=f(Y_{1}\mkern 1.0mu{|}\mkern 1.0muX)+\cdots+f(Y_{k}\mkern 1.0mu{|}\mkern 1.0muX),$
which similarly implies $f(Y_{I}\mkern 1.0mu{|}\mkern 1.0muX)=\sum_{i\in I}f(Y_{i}\mkern 1.0mu{|}\mkern 1.0muX)$ for every index set $I$.
@@ 3. The Maximum Entropy Method
Let us remark that while the maximum entropy extension is unique, the $m,n$-copy in Definition 1 is typically not, as the definition captures only a small part of the properties of the maximum entropy extension.
@@
$\mathcal{Q}=\big\{(\mathbf{x},\mathbf{u})\in\mathbb{R}^{d_{1}+d_{2}}:{}$ $B\mkern 1.5mu{\cdot}\mkern 1.5mu(\mathbf{x},\mathbf{u})^{T}\geqslant 0,\mbox{ and}$

### v2

@@ Abstract.
The entropic region is formed by the collection of the Shannon entropies of all subvectors of finitely many jointly distributed discrete random variables. For four or more variables, the structure of the entropic region is mostly unknown. We utilize a variant of the Maximum Entropy Method to obtain five-variable non-Shannon entropy inequalities, which delimit the five-variable entropy region. This method adds copies of some of the random variables in generations. A significant reduction in computational complexity, achieved through theoretical considerations and by harnessing the inherent symmetries, allowed us to calculate all five-variable non-Shannon inequalities provided by the first nine generations. Based on the results, we define two infinite collections of such inequalities and prove them to be entropy inequalities. We investigate downward-closed subsets of non-negative lattice points that parameterize these collections, and based on this, we develop an algorithm to enumerate all extremal inequalities. The discovered set of entropy inequalities is conjectured to characterize the applied method completely.
@@ 1. Introduction
Many important mathematical problems can be reduced to the following question: does a collection of finite random variables exist such that the entropies of the variable subsets satisfy certain linear constraints? Examples include, but are not limited to, channel coding [13] and network coding in particular [38], estimating the efficiency of secret sharing schemes [3, 4, 19], questions about matroid representations [2], guessing games [33, 34], extracting information from common strings in cryptography [17], additive combinatorics [23], and finding conditional independence inference rules [36].
@@
Previous works on generating and applying non-Shannon entropy inequalities, such as [4, 8, 9, 15, 29, 36], focused on the four-variable case, and only a few sporadic five-variable non-Shannon inequalities have been discovered, such as the MMRV inequality from [24] or an infinite set of inequalities from [28]. This is the first work that provides a method that systematically generates an infinite collection of non-Shannon bounds on the five-variable entropy region $\Gamma^{*}_{\!5}$. Compared to the four-variable case, there are significant challenges, both theoretical and computational. The four-variable entropy region $\Gamma^{*}_{\!4}$ sits in the $15$-dimensional Euclidean space, while the five-variable region $\Gamma^{*}_{\!5}$ is $31$-dimensional. The structure of the Shannon bound $\Gamma_{\!4}$ is well-understood: it has $41$ extremal directions, and only $6$ of them have no entropic points. The entropy region $\Gamma^{*}_{\!4}$ has an inner polyhedral cone where it fills its Shannon bound, and has six isomorphic “protrusions” towards the six exceptional extremal directions, each protrusion surrounded by $15$ hyperplanes of which $14$ come from the Shannon bound [25]. Only the protrusions contribute to new entropy inequalities, and their dimension can be reduced to $10$. Computational results about $\Gamma^{*}_{\!4}$ can be obtained by computing vertices and facets of numerous implicitly defined $10$-dimensional polyhedra [9]. In contrast, the Shannon bound $\Gamma_{\!5}$ of the five-variable entropy region has 117,983 extremal directions [35], and for a few of them it is not even known whether they contain an entropic point or not. No structural reduction similar to the four-variable case is available, and it is not known whether such a reduction exists or not. Computations about $\Gamma^{*}_{\!5}$ can still be reduced to $25$-dimensional polyhedral enumeration problems (although with significantly larger number of constraints than in the $4$-variable case). The complexity of enumeration problems typically doubles when the dimension increases by one, making such high-dimensional enumeration problems practically intractable.
@@
We overcome this computational difficulty by applying a particular variant of the Maximum Entropy Method. This variant, working in generations, first reduces the problem dimension from $31$ to $19$, and then, at each generation, adds extra copies of some of the random variables, increasing the problem dimension again. Theoretical considerations and harnessing the inherent symmetry allowed us to complete the associated polyhedral computations up to nine generations. The output was the complete list of five-variable non-Shannon inequalities provided by the first nine generations. Based on the experimental results, we define an infinite collection of five-variable inequalities that we prove to be provided by this MEM variant—in particular, they are valid non-Shannon entropy inequalities—and conjecture this collection to be complete; that is, no additional inequalities are yielded by this MEM variant. The collection of the inequalities is parametrized by finite, downward closed subsets of the non-negative lattice points of the plane. Some of the inequalities in our collection are consequences of the others; those that are not, are called extremal. We developed an incremental algorithm that enumerates, from generation to generation, the parameters yielding the extremal inequalities, in complete agreement with the computational results. The algorithm allowed us to significantly exceed the capabilities of polyhedral computation. While numerical instability prevented the completion of the polyhedral computation for the tenth generation, all extremal entropy inequalities were enumerated up to generation 60. Finally, we have looked at the large-scale behavior of the extremal inequalities, and pictured how these inequalities delimit a three-dimensional cross-section of $\Gamma^{*}_{\!5}$.
The new five-variable non-Shannon inequalities can be applied to real-world problems. The most immediate application is in network coding. The new inequalities tighten the boundaries; they provide stricter and more accurate bounds on network capacity. In a network protocol they can assist in proving whether a targeted data rate is achievable or not [38].
Cloud storage services (like Google Drive or AWS S3) distribute data fragments across many nodes [30]. In case of failure, the node has to download the missing data from other nodes. The new five-variable inequalities can be used to determine the theoretical limits of storage efficiency for systems with more complex failure models or larger clusters.
In the realm of secret sharing, entropy inequalities provide lower bounds on the size of secrets [3, 4, 19]. To explore another facet of these problems, the new inequalities can prove that certain efficient schemes are impossible to realize. In complex datasets, it is important to distinguish between correlation and actual causation [20]. When an AI model analyzes data to build a causal graph, it can use entropy inequalities to rule out models that are information-theoretically impossible, narrowing the search space and improving accuracy.
In this paper lemmas, claims and theorems are arranged so that each is used in the same section only, typically right after they are stated and proved. Section 4 proves structural properties of the entropy region that are used to reduce the computational complexity of the polyhedral algorithms. The main theoretical results are stated and proved in Section 7. In Theorem 11 we prove a large collection of entropy inequalities parametrized by the downward-closed subsets of the non-negative lattice points. Lemmas estimating different entropy expressions are used in this section only. Unfortunately, many inequalities provided by Theorem 11 are consequences of the others. Claims and lemmas in Section 8 provide the theoretical foundation for our algorithm that selects and enumerates the extremal inequalities among them.
@@
The remaining part of the paper is organized as follows. Notations are recalled in Section 2. Section 3 describes the special variant of the Maximum Entropy Method we apply to $\Gamma^{*}_{\!5}$. Section 4 discusses possible simplifications, including how symmetry can be utilized and how the MEM parameters were chosen. Section 5 describes the chosen coordinate systems, polyhedral computations, and their results. Section 6 presents the five-variable inequalities we obtained, paving the way for the definition of two infinite families of such inequalities in Section 7. Additional theoretical results, including the proof that inequalities in these families are indeed generated by the MEM method, are presented in Section 7. Section 8 discusses methods that can recognize extremal inequalities and the incremental algorithm that enumerates the extremal inequalities for each MEM generation, describes the large-scale behavior of the new inequalities, and investigates the delimited part of the five-variable entropy region. In Section 9, we take a look at the five-variable non-Shannon entropy inequalities constructed by F. Matúš in [28] and compare them to a special subclass of the inequalities obtained in Section 7. Finally, Section 10 summarizes our work, lists open questions, and provides directions for further research.
@@ 2. Preliminaries
A discrete random variable $\xi$ takes its values from a finite set $\mathcal{X}$, called alphabet. The probability that $\xi$ takes $x\in\mathcal{X}$ is denoted by $\Pr(\xi=x)$, or simply by $\Pr(x)$ when the random variable $\xi$ is clear from the context. Suppose $\xi$ is defined on the direct product $\mathcal{X}=\prod_{i\in N}\mathcal{X}_{i}$ for some finite set $N$, called the base set. For a non-empty $A\subseteq N$ the marginal $\xi_{A}$ is defined on the product alphabet $\mathcal{X}_{A}=\prod_{i\in A}\mathcal{X}_{i}$ so that the probability of $y\in\mathcal{X}_{A}$ is the sum of the probabilities of those $x\in\mathcal{X}$ whose projection to $\mathcal{X}_{A}$ equals $y$:
@@
with the convention that $0\log 0=0$. If $\xi=(\xi_{i}:i\in N)$ is a joint distribution, then we write $\mathbf{H}_{\xi}(A)$ for $\mathbf{H}(\xi_{A})$. The index $\xi$ is also dropped when it is clear from the context. By convention, $\mathbf{H}(\emptyset)=0$. The entropies $\mathbf{H}_{\xi}(A)$ are arranged into a vector indexed by the non-empty subsets $A$ of $N$. This vector is the entropy profile of the distribution $\xi$. The collection of these $(2^{|N|}\mkern 1.5mu{-}\mkern 1.5mu1)$-dimensional vectors forms the entropy region, denoted by $\Gamma^{*}_{\!N}$. Elements of $\Gamma^{*}_{\!N}$ are considered interchangeably as vectors, as points in this Euclidean space, and as functions assigning non-negative real numbers to non-empty subsets of the base set $N$. For a gentle introduction to these notions of Information Theory, please consult [37].
@@
$f(A\mkern 1.5mu{|}\mkern 1.5muB)$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}f(AB)-f(B),$
@@
$f(A,B\mkern 1.5mu{|}\mkern 1.5muC)$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}f(AC)+f(BC)-f(ABC)-f(C),\mbox{ and}$
$f[A,B,C,D]$ $\stackrel{{\scriptstyle\mathrm{def}}}{{=}}-f(A,B)+f(A,B\mkern 1.5mu{|}\mkern 1.5muC)+f(A,B\mkern 1.5mu{|}\mkern 1.5muD)+f(C,D).$
The first three expressions are called conditional entropy, mutual information, and conditional mutual information, respectively. The last line defines the Ingleton expression. An entropy function is not defined on the empty set, nevertheless, $f(\emptyset)=0$ will be assumed whenever convenient. In particular, $f(A,B\mkern 1.5mu{|}\mkern 1.5mu\emptyset)$ and $f(A,B)$ are the same expressions. Frequently, when clear from the context, the function $f$ is omitted before the parenthesized expression. Additionally, if applied to singletons, the Ingleton expression is written without commas. An example is the inequality
$abcd\mkern 1.5mu]+(a,b\mkern 1.5mu{|}\mkern 1.5muz)+(b,z\mkern 1.5mu{|}\mkern 1.5mua)+(a,z\mkern 1.5mu{|}\mkern 1.5mub)+3(z\mkern 1.5mu{|}\mkern 1.5muab)\geqslant 0.$ (1)
The Shannon inequalities state the non-negativity of the conditional entropy, mutual information, and conditional mutual information for all subsets $A$, $B$, $C$ of the base set $N$. They are consequences of the unique minimal set of such inequalities, called basic Shannon inequalities, see [37], listed in (B1) and (B2) below:
@@
    $f(i\mkern 1.5mu{|}\mkern 1.5muN\mkern 1.5mu{-}\mkern 1.5mui)\geqslant 0$ for all $i\in N$;
@@
    $f(a,b\mkern 1.5mu{|}\mkern 1.5muK)\geqslant 0$ for all $K\subseteq N$ and different $a,b\in N\mkern 1.5mu{-}\mkern 1.5muK$, including $K=\emptyset$.
The collection of all $(2^{|N|}\mkern 1.5mu{-}\mkern 1.5mu1)$-dimensional vectors (or points, or functions) that satisfy the Shannon inequalities is denoted by $\Gamma_{\!N}$. It is a natural outer bound for the entropy region $\Gamma^{*}_{\!N}$. $\Gamma_{\!N}$ is a pointed polyhedral cone [40]; its facets are the hyperplanes specified by the basic Shannon inequalities. Polymatroids are elements of $\Gamma_{\!N}$ written in functional form. A polymatroid is usually written as $(f,N)$, or just $f$, and we say that $f$ is on $N$. The polymatroid $f$ is entropic if it is in $\Gamma^{*}_{\!N}$, and almost entropic, or aent for short, if it is in the closure (in the usual Euclidean topology) of $\Gamma^{*}_{\!N}$. Linear inequalities valid for all polymatroids are consequences of the basic Shannon inequalities; an example is the inequality (1). A non-Shannon inequality is a homogeneous linear inequality that is valid for points of the entropic region but not for all points of the Shannon bound. Equivalently, the non-negative side of the hyperplane corresponding to such an inequality contains the complete entropy region, while it cuts properly into $\Gamma_{\!N}$.
@@
In this case we also have $f(Y_{I})=\sum_{i\in I}f(Y_{i})$ for every subset $I$ of the indices. The disjoint subsets $Y_{1}$ and $Y_{2}$ are conditionally independent over $X$ if $f(Y_{1},Y_{2}\mkern 1.5mu{|}\mkern 1.5muX)=0$; and $Y_{1},\dots,Y_{k}$ are completely conditionally independent over $X$ if $Y_{I}$ and $Y_{J}$ are conditionally independent over $X$ for arbitrary disjoint subsets $I$ and $J$ of the indices. An equivalent condition is
$f(Y_{1}\cdots Y_{k}\mkern 1.5mu{|}\mkern 1.5muX)=f(Y_{1}\mkern 1.5mu{|}\mkern 1.5muX)+\cdots+f(Y_{k}\mkern 1.5mu{|}\mkern 1.5muX),$
which similarly implies $f(Y_{I}\mkern 1.5mu{|}\mkern 1.5muX)=\sum_{i\in I}f(Y_{i}\mkern 1.5mu{|}\mkern 1.5muX)$ for every index set $I$.
@@ 3. The Maximum Entropy Method
Let us remark that while the maximum entropy extension is unique, the $m,n$-copy in Definition 1 is typically not, as the definition captures only a small part of the properties of the maximum entropy extension. The obtained entropy inequalities form the facets of a convex polytope; consequently, they are independent in the sense that none of them is a consequence of the others or the Shannon inequalities.
@@
$\mathcal{Q}=\big\{(\mathbf{x},\mathbf{u})\in\mathbb{R}^{d_{1}+d_{2}}:{}$ $B\mkern 1.5mu{\cdot}\mkern 1.5mu(\mathbf{x},\mathbf{u})^{\top}\geqslant 0,\mbox{ and}$